jakfishman:

If y = 6x − 3, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? (4 points) {(0, 3), (1, −9), (2, −15)} {(0, −3), (2, 15), (4, 21)} {(0, −3), (2, 9), (4, 21)} {(0, 3), (−2, 15), (4, 21)} plz help

mxddi3:

ok so since you have ordered pairs, just plug the numbers in for x and y. like for the first one, your first ordered pair is (0,3) so make y=3 and x=0 \[3=6(0)-3\] simplify and get 3=0-3----> 3=-3 since this isn't true, we can eliminate that answer choice. now can you do the same thing with the next set please?

jakfishman:

thx

jakfishman:

if im corect its number 3 right

mxddi3:

if the first ordered pair is correct, be sure to check the second and third ones because they could be incorrect making it not true

mxddi3:

yeah

jakfishman:

ok thank you